1. Chromium (Cr) is a metal that is added to steel to improve its resistance to corrosion (for example, to make stainless steel).  Calculate both the number of moles in a sample of chromium containing 5.00 x 1020 atoms and the mass of the sample.
1
Expert's answer
2021-05-06T07:39:20-0400

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

(5.00×1020 atoms of Cr) × (1 mol / 6.022×1023 atoms) = 0.00083 mol = 8.30×10-4 mol Cr

Moles of Cr = 8.30×10-4 mol


The molar mass of chromium (Cr) is 51.9961 g/mol.

Hence,

(8.30×10-4 mol Cr) × (51.9961 g Cr / 1 mol Cr) = 0.043 g Cr

Mass of Cr = 0.043 g


Answers:

1) 8.30×10-4 mol Cr;

2) 0.043 g Cr.